4. Let A 7 3 3 -1 1 Find matrix P such that D = P-1AP is diagonal.
Answer #1

binen that As 3-) fou eigenvalu [A – AI) = 0 3-) 3 3 --) (3-») (-1-я) -9 = 0 -1 -1, +, +7? – go 22 5 - 16 = 0 Х~ 8, +2-10 - 0